# Higher-Rank Irreducible Cartesian Tensors for Equivariant Message Passing

### v1

@@ Abstract
The ability to perform fast and accurate atomistic simulations is crucial for advancing the chemical sciences. By learning from high-quality data, machine-learned interatomic potentials achieve accuracy on par with ab initio and first-principles methods at a fraction of their computational cost. The success of machine-learned interatomic potentials arises from integrating inductive biases such as equivariance to group actions on an atomic system, e.g., equivariance to rotations and reflections. In particular, the field has notably advanced with the emergence of equivariant message-passing architectures. Most of these models represent an atomic system using spherical tensors, tensor products of which require complicated numerical coefficients and can be computationally demanding. This work introduces higher-rank irreducible Cartesian tensors as an alternative to spherical tensors, addressing the above limitations. We integrate irreducible Cartesian tensor products into message-passing neural networks and prove the equivariance of the resulting layers. Through empirical evaluations on various benchmark data sets, we consistently observe on-par or better performance than that of state-of-the-art spherical models.
@@ 1 Introduction
The ability to perform sufficiently fast and accurate atomistic simulations for large molecular and material systems holds the potential to revolutionize molecular and materials science [1, 2, 3, 4, 5, 6, 7, 8]. Conventionally, computational chemistry and materials science rely on ab initio or first-principles approaches—e.g., coupled cluster [9, 10, 11] or density functional theory (DFT) [12, 13], respectively—that are accurate but computationally demanding, thus limiting the accessible simulation time and system sizes. However, the ability to generate high-quality, first-principles-based data sets has prompted the development of machine-learned interatomic potentials (MLIPs). These potentials enable atomistic simulations with accuracy that is on par with the reference first-principles method but at a fraction of the computational cost. Message-passing neural networks (MPNNs) [14, 15, 16, 17, 18, 19, 20, 21] have been employed in chemical and materials sciences, including the development of MLIPs, due to their efficient processing of the graph representation of the atomic system [22, 23, 24, 25, 26]. Achieving the desired MLIPs’ performance, however, requires the inclusion of inductive biases, like the invariance of total energy to translations, reflections, and rotations in the three-dimensional space, and an effective encoding of the atomic system into a learnable representation [27]. Designing equivariant MPNNs [28, 29, 30, 31, 32, 25, 33, 34, 35, 36, 26], which preserve the directional information of the local atomic environment, has been one of the most active research directions of the last years. Previous work often achieves equivariance by representing an atomic system as a graph where node features are expressed in spherical harmonics basis and designing ad-hoc message-passing layers [25, 32, 33, 35, 34, 36].
@@
MLIPs based on spherical harmonics [25, 33, 34], which are the basis for irreducible representations of the three-dimensional rotation group, often demonstrate a better performance compared to those that use lower-rank Cartesian representations (scalars and vectors) [28, 37]. Spherical tensors, however, require the definition of a particular rotational axis, often chosen as the $z$-axis, resulting in inherent bias [38]. Furthermore, coupling spherical tensors via tensor products, involved in designing equivariant convolutions and constructing many-body features [39, 40, 33], requires the definition of complicated numerical coefficients, e.g., Wigner $3\mbox{-}j$ symbols defined in terms of the Clebsch–Gordan coefficients [41], and can be computationally demanding. In contrast, irreducible Cartesian tensors have no preferential directions; their tensor products are simpler and have a better computational complexity—up to a certain tensor rank—than the tensor products of spherical tensors [38, 42, 43, 44, 45]. Recent work [30] improved results of Cartesian MPNNs by decomposing rank-two tensors into irreducible representations of the three-dimensional rotation group, i.e., into representations of dimension 1 (trace), 3 (anti-symmetric part), and 5 (traceless symmetric part) [46].
@@
Contributions. In this work, we demonstrate that operating with irreducible Cartesian tensors leads to on-par or, sometimes, even better performance than that of spherical counterparts. Particularly, this work goes beyond scalars, vectors, and rank-two tensors and explores the integration of higher-rank irreducible Cartesian tensors and their products into equivariant MPNNs: i) We demonstrate how irreducible Cartesian tensors that are symmetric and traceless can be constructed from a unit vector and a product of two irreducible Cartesian tensors; ii) We prove that the resulting tensors are equivariant under the action of the three-dimensional rotation and reflection group; iii) We demonstrate that higher-rank irreducible Cartesian tensors can be used to design cost-efficient—up a certain tensor rank—and accurate equivariant models of many-body interactions; iv) We conduct different experiments to assess the effectiveness of equivariant message passing based on irreducible Cartesian tensors, and achieve state-of-the-art performance on benchmark data sets such as rMD17 [47], 3BPA [48], and acetylacetone [32]. We hope our contributions will offer new insights into the use of Cartesian tensors for constructing accurate and cost-efficient MLIPs and beyond.
@@ 2 Background
Cartesian tensors. A vector $\mathbf{x}\in\mathbb{R}^{3}$ transforms under the action of the rotation group $\mathrm{SO}(3)$ as $\mathbf{x}^{\prime}=R\mathbf{x}$, i.e., each component of it transforms as $\left(\mathbf{x}\right)_{i}^{\prime}=\sum_{j}R_{ij}\left(\mathbf{x}\right)_{j}$. Here, $R=D_{\mathcal{X}}\left[g\right]\in\mathbb{R}^{3\times 3}$ denotes the rotation matrix representation of $g\in\mathrm{SO}(3)$. Cartesian tensors of rank $n$ are described by $3^{n}$ numbers and generalize the concept of vectors. A rank-$n$ Cartesian tensor $\mathbf{T}$ can be viewed as a multidimensional array with $n$ indices, i.e., $\left(\mathbf{T}\right)_{i_{1}i_{2}\cdots i_{n}}$ with $i_{k}\in\{1,2,3\}$ for $\forall\,k\in\{1,\cdots,n\}$. Furthermore, each index of $\left(\mathbf{T}\right)_{i_{1}i_{2}\cdots i_{n}}$ transforms independently as a vector under rotation. For example, a rank-two tensor transforms under rotation as $\mathbf{T}^{\prime}=R\mathbf{T}R^{\top}$, i.e., each component of it transforms as $\left(\mathbf{T}\right)_{i_{1}i_{2}}^{\prime}=\sum_{j_{1}j_{2}}R_{i_{1}j_{1}}R_{i_{2}j_{2}}\left(\mathbf{T}\right)_{j_{1}j_{2}}$. For Cartesian tensors, one defines an $r$-fold tensor contraction and an outer product, i.e., $\left(\mathbf{T}\right)_{j_{1}\cdots j_{s}}=\sum_{i_{1},\cdots,i_{r}}\left(\mathbf{U}\right)_{i_{1}\cdots i_{r}j_{1}\cdots j_{s}}\left(\mathbf{S}\right)_{i_{1}\cdots i_{r}}$ and $\left(\mathbf{T}\right)_{i_{1}\cdots i_{r}j_{1}\cdots j_{s}}=\left(\mathbf{U}\right)_{i_{1}\cdots i_{r}}\left(\mathbf{S}\right)_{j_{1}\cdots j_{s}}$, respectively. Cartesian tensors are generally reducible and can, thus, be decomposed into smaller representations that transform independently within their linear subspaces under rotation. For example, a rank-two reducible Cartesian tensor contains representations of dimension 1 (trace), 3 (anti-symmetric part), and 5 (traceless symmetric part): $\sum_{i_{1}}(\mathbf{T})_{i_{1}i_{1}}$, $(\mathbf{T})_{i_{1}i_{2}}-(\mathbf{T})_{i_{2}i_{1}}$, and $(\mathbf{T})_{i_{1}i_{2}}+(\mathbf{T})_{i_{2}i_{1}}-2/3\sum_{i_{1}}(\mathbf{T})_{i_{1}i_{1}}$, respectively. Under rotation, the symmetric traceless part remains within its irreducible subspace, with a similar behavior for other irreducible components of $(\mathbf{T})_{i_{1}i_{2}}$. Furthermore, for a reducible Cartesian tensor of rank $n$, an action of the rotation group $\mathrm{SO}(3)$ can be represented with a $3^{n}\times 3^{n}$-dimensional rotation matrix. This rotation matrix is also reducible, meaning that an appropriate change of basis can block diagonalize it into smaller subsets, the irreducible representations of the rotation group $\mathrm{SO}(3)$.
@@ 3 Related work
Higher-body-order local descriptors. Recent advances in MLIPs have been influenced by moment tensor potentials (MTP) [50] and atomic cluster expansion (ACE) [40]. These approaches enable systematic construction of higher-body-order polynomial basis functions, encompassing representations like atom-centered symmetry functions (ACSFs) [51, 52], smooth overlap of atomic positions (SOAP) [53, 54], Gaussian moments [55, 56], and embedded atom/multi-scale embedded atom method (EAM/MEAM) potentials [57, 58]. Furthermore, a reducible Cartesian tensor can be represented as a linear combination of irreducible spherical counterparts, and vice versa [40, 49]. More general expressions, including tensor contractions, have also been provided, hinting at the relationship between Cartesian and spherical models [59, 60, 61]. Despite the success of MTP and ACE, defining smaller cutoff radii and rigid architecture can result in limited accuracy compared to MPNNs.
@@
Equivariant message-passing potentials. Equivariant MPNNs [39, 62, 28, 63, 64, 25, 37, 32, 33, 29, 35, 34, 30, 36, 31] often outperform more traditional invariant counterparts [65, 66, 67, 68, 69, 70]. While many invariant and equivariant MPNNs rely on two-body features within a single message-passing layer, there is a growing interest in incorporating higher-body-order features, such as angles and dihedrals, to model many-body interactions in atomic systems [69, 70, 67]. Note that equivariant models with two-body features in a single message-passing layer build many-body ones through numerous message-passing iterations, increasing the receptive field and, thus, computational cost. Refs. 67, 69 recognized the importance of higher-body-order features but faced challenges due to explicit summation over triplets or quadruplets. In contrast, MACE advances the current state-of-the-art by combining ACE and equivariant message passing, introducing cost-efficient many-body message-passing potentials [33]. With just two message-passing layers, MACE yields accurate potentials for interacting many-body systems [71].
@@
Beyond Clebsch–Gordan tensor product. Despite the success of equivariant MPNNs based on spherical tensors, the high computational cost of the Clebsch–Gordan tensor product limits their computational efficiency [72, 28, 63, 73, 33, 64, 34, 35, 30, 74]. Thus, current research focuses on alternatives to the Clebsch–Gordan tensor product, which has a $\mathcal{O}(L^{6})$ complexity for tensors up to degree $L$. Recently, the relation of Clebsch–Gordan coefficients to the integral of products of three spherical harmonics, known as the Gaunt coefficients [41], has been exploited to reduce the computational cost of the tensor product of spherical tensors to $\mathcal{O}(L^{3})$ [36]. Alternatively, Cartesian tensors and their products can be used to construct message-passing layers with two- and many-body features. Recently, a competitive performance to spherical models for a Cartesian message-passing architecture based on decomposing rank-two tensors into their irreducible representations has been demonstrated [30]. In this case, all operations are performed on tensors or their decomposition, which has advantageous computational complexity compared to the Clebsch–Gordan tensor product. Furthermore, recent work employs higher-rank reducible Cartesian tensors to design many-body message-passing potentials but does not achieve the state-of-the-art performance of spherical counterparts [31]. This performance gap may result from mixing different irreducible representations during massage-passing based on reducible Cartesian tensors and their products.
@@ 4.1 Irreducible Cartesian tensor product
Figure 1: Schematic illustration of (a) the construction of an irreducible Cartesian tensor for a local atomic environment and (b) the tensor product of two irreducible Cartesian tensors of rank $l_{1}$ and $l_{2}$. The construction of an irreducible Cartesian tensor from a unit vector $\hat{\mathbf{r}}$ is defined in Eq. (1). In this work, we use tensors with the same rank $n$ and weight $l$, i.e., $n=l$, avoiding the need for embedding tensors with $l<n$ in a higher-dimensional tensor space. Therefore, we use $l$ to identify the rank and the weight of an irreducible Cartesian tensor throughout this work. The tensor product is defined in Eqs. (2) and (3), resulting in a new tensor $\mathbf{T}_{l_{3}}=(\mathbf{T}_{l_{1}}\otimes_{\mathrm{Cart}}\mathbf{T}_{l_{2}})_{l_{3}}$ of rank $l_{3}=\{\lvert l_{1}-l_{2}\rvert,\cdots,l_{1}+l_{2}\}$. Transparent boxes denote the linearly dependent elements of symmetric and traceless tensors. The tensor product yields even and odd parity tensors $\mathbf{T}_{l_{3}}$, given by $l_{1}+l_{2}-l_{3}$.
@@
$\begin{split}(\mathbf{x}_{l_{1}}&\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}=\\ &C_{l_{1}l_{2}l_{3}}\sum\nolimits_{m=0}^{\mathrm{min}(l_{1},l_{2})-k}(-1)^{m}2^{m}\frac{(2l_{3}-2m-1)!!}{(2l_{3}-1)!!}\big{\{}\left(\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)\otimes\mathbf{I}^{\otimes m}\big{\}},\end{split}$ (2)
@@
$\begin{split}(\mathbf{x}_{l_{1}}&\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}=\\ &D_{l_{1}l_{2}l_{3}}\sum\nolimits_{m=0}^{\mathrm{min}(l_{1},l_{2})-k-1}(-1)^{m}2^{m}\frac{(2l_{3}-2m-1)!!}{(2l_{3}-1)!!}\big{\{}\left(\boldsymbol{\varepsilon}:\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)\otimes\mathbf{I}^{\otimes m}\big{\}},\end{split}$ (3)

### v2

@@ Abstract
The ability to perform fast and accurate atomistic simulations is crucial for advancing the chemical sciences. By learning from high-quality data, machine-learned interatomic potentials achieve accuracy on par with ab initio and first-principles methods at a fraction of their computational cost. The success of machine-learned interatomic potentials arises from integrating inductive biases such as equivariance to group actions on an atomic system, e.g., equivariance to rotations and reflections. In particular, the field has notably advanced with the emergence of equivariant message passing. Most of these models represent an atomic system using spherical tensors, tensor products of which require complicated numerical coefficients and can be computationally demanding. Cartesian tensors offer a promising alternative, though state-of-the-art methods lack flexibility in message-passing mechanisms, restricting their architectures and expressive power. This work explores higher-rank irreducible Cartesian tensors to address these limitations. We integrate irreducible Cartesian tensor products into message-passing neural networks and prove the equivariance and traceless property of the resulting layers. Through empirical evaluations on various benchmark data sets, we consistently observe on-par or better performance than that of state-of-the-art spherical and Cartesian models.
@@ 1 Introduction
11footnotetext: Corresponding author: viktor.zaverkin@neclab.eu22footnotetext: These authors contributed equally.
@@
The ability to perform sufficiently fast and accurate atomistic simulations for large molecular and material systems holds the potential to revolutionize molecular and materials science [1, 2, 3, 4, 5, 6, 7, 8]. Conventionally, computational chemistry and materials science rely on ab initio or first-principles approaches—e.g., coupled cluster [9, 10, 11] or density functional theory (DFT) [12, 13], respectively—that are accurate but computationally demanding, thus limiting the accessible simulation time and system sizes. However, the ability to generate high-quality, first-principles-based data sets has prompted the development of machine-learned interatomic potentials (MLIPs). These potentials enable atomistic simulations with accuracy that is on par with the reference first-principles method but at a fraction of the computational cost. Message-passing neural networks (MPNNs) [14, 15, 16, 17, 18, 19, 20, 21] have been employed in chemical and materials sciences, including the development of MLIPs, due to their efficient processing of the graph representation of the atomic system [22, 23, 24, 25, 26]. Achieving the desired MLIPs’ performance, however, requires the inclusion of inductive biases, like the invariance of total energy to translations, reflections, and rotations in the three-dimensional space, and an effective encoding of the atomic system into a learnable representation [27]. Designing equivariant MPNNs [28, 29, 30, 31, 32, 25, 33, 34, 35, 36, 26], which preserve the directional information of the local atomic environment, has been one of the most active research directions of the last years. Previous work often achieves equivariance by representing an atomic system as a graph, with node features expressed in spherical harmonics basis or using lower-rank Cartesian tensors and designing ad-hoc message-passing layers [28, 29, 25, 32, 33, 37, 30, 35, 34, 36, 31].
@@
MLIPs based on spherical harmonics [25, 33, 34], which are the basis for irreducible representations of the three-dimensional rotation group, often demonstrate a better performance compared to those that use lower-rank Cartesian representations (scalars and vectors) [28, 37]. Spherical tensors, however, require the definition of a particular rotational axis, often chosen as the $z$-axis, resulting in inherent bias [38]. Furthermore, coupling spherical tensors via tensor products, involved in designing equivariant convolutions and constructing many-body features [39, 40, 33], requires the definition of complicated numerical coefficients, e.g., Wigner $3\mbox{-}j$ symbols defined in terms of the Clebsch–Gordan coefficients [41], and can be computationally demanding. In contrast, irreducible Cartesian tensors have no preferential directions; their tensor products are simpler and have a better computational complexity—up to a certain tensor rank—than the tensor products of spherical tensors [38, 42, 43, 44, 45]. Recent work improved results of Cartesian MPNNs by using rank-two tensors and decomposing them into irreducible representations of the three-dimensional rotation group [30], i.e., into representations of dimension 1 (trace), 3 (anti-symmetric part), and 5 (traceless symmetric part) [46]. Higher-rank reducible Cartesian tensors have also been integrated into many-body MPNNs [31]. These state-of-the-art Cartesian approaches, however, lack the flexibility of their message-passing mechanisms. They rely exclusively on convolutions with invariant filters and restrict the construction of many-body features, limiting the range of possible architectures and their expressive power.
@@
Contributions. In this work, we address the limitations of state-of-the-art Cartesian models and demonstrate that operating with irreducible Cartesian tensors leads to on-par or, sometimes, even better performance than that of spherical counterparts. Particularly, this work goes beyond scalars, vectors, and rank-two tensors and explores the integration of higher-rank irreducible Cartesian tensors and their products into equivariant MPNNs: i) We demonstrate how irreducible Cartesian tensors that are symmetric and traceless can be constructed from a unit vector and a product of two irreducible Cartesian tensors, essential for equivariant convolutions and constructing many-body features; ii) We prove that the resulting tensors are traceless and equivariant under the action of the three-dimensional rotation and reflection group; iii) We demonstrate that higher-rank irreducible Cartesian tensors can be used to design cost-efficient—up to a certain tensor rank—and accurate equivariant models of many-body interactions; iv) We conduct different experiments to assess the effectiveness of equivariant message passing based on irreducible Cartesian tensors, and achieve state-of-the-art performance on benchmark data sets such as rMD17 [47], MD22 [48], 3BPA [49], acetylacetone [32], and Ta–V–Cr–W [50]. We hope our contributions will offer new insights into the use of Cartesian tensors for constructing accurate and cost-efficient MLIPs and beyond.
@@ 2 Background
Cartesian tensors. A vector $\mathbf{x}\in\mathbb{R}^{3}$ transforms under the action of the rotation group $\mathrm{SO}(3)$ as $\mathbf{x}^{\prime}=R\mathbf{x}$, i.e., each component of it transforms as $\left(\mathbf{x}\right)_{i}^{\prime}=\sum_{j}R_{ij}\left(\mathbf{x}\right)_{j}$. Here, $R=D_{\mathcal{X}}\left[g\right]\in\mathbb{R}^{3\times 3}$ denotes the rotation matrix representation of $g\in\mathrm{SO}(3)$. Cartesian tensors of rank $n$ are described by $3^{n}$ numbers and generalize the concept of vectors. A rank-$n$ Cartesian tensor $\mathbf{T}$ can be viewed as a multidimensional array with $n$ indices, i.e., $\left(\mathbf{T}\right)_{i_{1}i_{2}\cdots i_{n}}$ with $i_{k}\in\{1,2,3\}$ for $\forall\,k\in\{1,\cdots,n\}$. Furthermore, each index of $\left(\mathbf{T}\right)_{i_{1}i_{2}\cdots i_{n}}$ transforms independently as a vector under rotation. For example, a rank-two tensor transforms under rotation as $\mathbf{T}^{\prime}=R\mathbf{T}R^{\top}$, i.e., each component of it transforms as $\left(\mathbf{T}\right)_{i_{1}i_{2}}^{\prime}=\sum_{j_{1}j_{2}}R_{i_{1}j_{1}}R_{i_{2}j_{2}}\left(\mathbf{T}\right)_{j_{1}j_{2}}$. For Cartesian tensors, one defines an $r$-fold tensor contraction and an outer product, i.e., $\left(\mathbf{T}\right)_{j_{1}\cdots j_{s}}=\sum_{i_{1},\cdots,i_{r}}\left(\mathbf{U}\right)_{i_{1}\cdots i_{r}j_{1}\cdots j_{s}}\left(\mathbf{S}\right)_{i_{1}\cdots i_{r}}$ and $\left(\mathbf{T}\right)_{i_{1}\cdots i_{r}j_{1}\cdots j_{s}}=\left(\mathbf{U}\right)_{i_{1}\cdots i_{r}}\left(\mathbf{S}\right)_{j_{1}\cdots j_{s}}$, respectively. Cartesian tensors are generally reducible and can, thus, be decomposed into smaller representations that transform independently within their linear subspaces under rotation. For example, a rank-two reducible Cartesian tensor contains representations of dimension 1 (trace), 3 (anti-symmetric part), and 5 (traceless symmetric part): $\sum_{i_{1}}(\mathbf{T})_{i_{1}i_{1}}$, $(\mathbf{T})_{i_{1}i_{2}}-(\mathbf{T})_{i_{2}i_{1}}$, and $(\mathbf{T})_{i_{1}i_{2}}+(\mathbf{T})_{i_{2}i_{1}}-2/3\sum_{i_{1}}(\mathbf{T})_{i_{1}i_{1}}$, respectively. Under rotation, the traceless symmetric part remains within its irreducible subspace, with a similar behavior for other irreducible components of $(\mathbf{T})_{i_{1}i_{2}}$. Furthermore, for a reducible Cartesian tensor of rank $n$, an action of the rotation group $\mathrm{SO}(3)$ can be represented with a $3^{n}\times 3^{n}$-dimensional rotation matrix. This rotation matrix is also reducible, meaning that an appropriate change of basis can block diagonalize it into smaller subsets, the irreducible representations of the rotation group $\mathrm{SO}(3)$.
@@ 3 Related work
Equivariant message-passing potentials. Equivariant MPNNs [39, 52, 28, 53, 54, 25, 37, 32, 33, 29, 35, 34, 30, 36, 31] often outperform more traditional invariant models [55, 56, 57, 58, 59, 60]. While many invariant and equivariant MPNNs rely on two-body features within a single message-passing layer, there is a growing interest in constructing higher-body-order features, such as angles and dihedrals, to model many-body interactions in atomic systems [59, 60, 57]. Note that equivariant models with two-body features in a single message-passing layer build many-body ones through repeated message-passing iterations, increasing the receptive field and, thus, computational cost. Recent work recognized the importance of higher-body-order features but faced challenges due to explicit summation over triplets or quadruplets [57, 59]. In contrast, MACE advances the current state-of-the-art by combining ACE and equivariant message passing, introducing cost-efficient many-body message-passing potentials [33]. With just two message-passing layers, MACE yields accurate potentials for interacting many-body systems [61].
@@
Beyond Clebsch–Gordan tensor product. Despite the success of equivariant MPNNs based on spherical tensors, the high computational cost of the Clebsch–Gordan tensor product limits their computational efficiency [62, 28, 53, 63, 33, 54, 34, 35, 30, 64]. Thus, current research focuses on alternatives to the Clebsch–Gordan tensor product, which has a $\mathcal{O}(L^{5})$ complexity for tensors up to degree $L$. Recently, the relation of Clebsch–Gordan coefficients to the integral of products of three spherical harmonics, known as the Gaunt coefficients [41], has been exploited to reduce the computational cost of the tensor product of spherical tensors to $\mathcal{O}(L^{3})$, compared to $\mathcal{O}(L^{6})$ of the full Clebsch–Gordan tensor product including all $(l_{1},l_{2})\rightarrow l_{3}$ [36]. However, this approach excludes odd tensor products and, thus, restricts the expressive power and the range of possible architectures.
Cartesian tensors and their products offer another promising alternative to the Clebsch–Gordan tensor product, enabling the efficient construction of message-passing layers with two- and many-body features. Recent work has explored decomposing rank-two tensors into their irreducible representations and using higher-rank reducible tensors [30, 31]. These approaches, however, are limited to convolutions with invariant filters, restricting the range of possible architectures and, thus, the expressive power of resulting Cartesian models. Furthermore, they provide limited mechanisms for constructing higher-body-order features, restricted to three-body features obtained through matrix-matrix multiplication and invariant many-body features obtained through full tensor contractions, similar to moment tensor potentials and Gaussian moments [65, 66, 67]. Finally, using reducible Cartesian tensors during message-passing leads to mixing different irreducible representations, which can hinder the performance of resulting models compared to state-of-the-art spherical models [31].
@@ 4.1 Irreducible Cartesian tensor product
Figure 1: Schematic illustration of (a) the construction of an irreducible Cartesian tensor for a local atomic environment and (b) the tensor product of two irreducible Cartesian tensors of rank $l_{1}$ and $l_{2}$. The construction of an irreducible Cartesian tensor from a unit vector $\hat{\mathbf{r}}$ is defined in Eq. (1). In this work, we use tensors with the same rank $n$ and weight $l$, i.e., $n=l$, avoiding the need for embedding tensors with $l<n$ in a higher-dimensional tensor space. Therefore, we use $l$ to identify the rank and the weight of an irreducible Cartesian tensor. The tensor product is defined in Eqs. (2) and (3), resulting in a new tensor $\mathbf{T}_{l_{3}}=(\mathbf{T}_{l_{1}}\otimes_{\mathrm{Cart}}\mathbf{T}_{l_{2}})_{l_{3}}$ of rank $l_{3}=\{\lvert l_{1}-l_{2}\rvert,\cdots,l_{1}+l_{2}\}$. Transparent boxes denote the linearly dependent elements of symmetric and traceless tensors. The tensor product can be even or odd, defined by $l_{1}+l_{2}-l_{3}$.
@@
$\begin{split}(\mathbf{x}_{l_{1}}&\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}\\ =&C_{l_{1}l_{2}l_{3}}\sum\nolimits_{m=0}^{\mathrm{min}(l_{1},l_{2})-k}(-1)^{m}2^{m}\frac{(2l_{3}-2m-1)!!}{(2l_{3}-1)!!}\big{\{}\left(\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)\otimes\mathbf{I}^{\otimes m}\big{\}},\end{split}$ (2)
@@
$\begin{split}(\mathbf{x}_{l_{1}}&\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}\\ =&D_{l_{1}l_{2}l_{3}}\sum\nolimits_{m=0}^{\mathrm{min}(l_{1},l_{2})-k-1}(-1)^{m}2^{m}\frac{(2l_{3}-2m-1)!!}{(2l_{3}-1)!!}\big{\{}\left(\boldsymbol{\varepsilon}:\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)\otimes\mathbf{I}^{\otimes m}\big{\}},\end{split}$ (3)